# Inverted-circuit zero-noise extrapolation for quantum gate error mitigation

### v1

@@ Abstract
A common approach to deal with gate errors in modern quantum-computing hardware is zero-noise extrapolation. By artificially amplifying errors and extrapolating the expectation values obtained with different error strengths towards the zero-error (zero-noise) limit, the technique aims at rectifying errors in noisy quantum computing systems. For an accurate extrapolation, it is essential to know the exact factors of the noise amplification. In this article, we propose a simple method for estimating the strength of errors occurring in a quantum circuit and demonstrate improved extrapolation results. The method determines the error strength for a circuit by appending to it the inverted circuit and measuring the probability of the initial state. The estimation of error strengths is easy to implement for arbitrary circuits and does not require a previous characterisation of noise properties. We compare this method with the conventional zero-noise extrapolation method and show that the novel method leads to a more accurate calculation of expectation values. Our method proves to be particularly effective on current hardware, showcasing its suitability for near-term quantum computing applications.
error mitigation, zero-noise extrapolation
@@ I Introduction
Current quantum computing devices, so called Noisy Intermediate Scale Quantum (NISQ) hardware [1], suffer from errors [1, 2, 3, 4]. The accuracy of the output is reduced by these errors. Quantum error mitigation techniques aim to minimise these inaccuracies. The general principle of these techniques is to deduce an almost noise-free estimate of the expectation value of interest from results obtained on noisy hardware. For example, the technique of zero-noise extrapolation (ZNE) [2, 5, 6, 3, 7, 8] consists of performing a series of measurements with systematically amplified errors. The expectation values obtained with different error strengths are then extrapolated to zero error (zero-noise). Hence, a more accurate estimation of the expectation value can be obtained at the cost of an increased sampling overhead due to a larger number of circuits and classical post-processing.
@@
Several zero noise extrapolation schemes have been proposed, each with its advantages and challenges. In continuous ZNE [2, 5], the microwave pulses generating the elementary single- and two-qubit gates are scaled to increase the noise throughout the circuit. This extrapolation yields satisfactory results, however, it requires a time-consuming calibration of the scaled pulses. The implementation of the discrete ZNE method [9, 10, 11, 12] is more straightforward. Here, sequences of gates are inserted that, in the error-free case, are equivalent to the identity. Depending on the details of the implementation, different noise scaling factors can be targeted; for example, replacing every CNOT gate by three CNOT gates corresponds to a noise-scaling factor of three. While this simple scaling behaviour holds, for example, for depolarising errors in the limit of small error strength, this is usually not the case for realistic noise models. Depending on the specific type of errors occurring on the given hardware, not every error is amplified exactly $\lambda$ times if the corresponding gate is executed $\lambda$ times [13, 14, 15]. This may lead to a certain bias concerning the estimation of the error strength, which deteriorates the results of the discrete ZNE method [13, 14]. Another powerful approach is ZNE with Probabilistic Error Amplification (PEA) [16]. With PEA, the noise model is first learned for different circuit layers, and the errors are then amplified by inserting random gates according to the noise model. This approach has a high sampling cost due to the larger number of experiments required to learn the noise model. Probabilistic Error Cancellation (PEC) [2, 17] is another method with an even higher sampling cost [18] that represents the ideal circuit as a quasi-probability distribution of previously characterised noisy ones.
@@
In this article, by building on the relatively simple method of ZNE with identity insertions [9], we introduce the method of Inverted-Circuit-ZNE (IC-ZNE) which allows to measure the entire error of a circuit to obtain an exact noise scaling parameter for better error mitigation. This involves adding an extra circuit with its inverse to determine the actual error strength $\epsilon$, instead of measuring and mitigating the errors partially [19, 16, 2, 17]. The novel method overcomes the above problem that not every error type is amplified $\lambda$-times when increasing the number of noisy gates by $\lambda$ [6]. We further explicitly show that the total error is actually different from the assumed error amplification. The novel approach IC-ZNE does not require any knowledge of the underlying quantum computer noise model.
@@
The paper is organised as follows. After introducing known concepts and methods used in this paper like standard ZNE (sZNE) and randomised compiling in Sec. II, we introduce the method of Inverted-Circuit ZNE (IC-ZNE) in Sec. III. Starting from a precise definition of the error strength $\epsilon$, we derive a relation between $\epsilon$ and the probability to measure all qubits in their initial state after the inverse circuit has been added. Finally, we demonstrate and explain the improved performance of our method compared to standard ZNE on simulators with depolarising noise and on IBMs quantum computing hardware in Sec. IV. A conclusion is drawn in Sec. V.
@@ II.1 General setting and notation
The main idea of ZNE now consists of introducing a parameter $\epsilon$ which is expected to describe the amount of noise present in ${\mathcal{E}}_{U}$. Assuming that this parameter, or at least its scaling $\lambda=\epsilon/\epsilon_{0}$ with respect to the (in general unknown) noise strength $\epsilon_{0}$ of the original circuit, can be controlled by adding additional noise to the system, the data points $\langle A\rangle_{\lambda_{i}}$ measured at different noise scaling factors $\lambda_{i}$ can be extrapolated to yield $\langle A\rangle_{\rm ideal}=\langle A\rangle_{\lambda\to 0}$. Different versions of ZNE differ in the way how the noise parameter $\epsilon$ (or, equivalently, the noise scaling factor $\lambda$) is defined, the number of data points to be measured at different values of the noise strengths and the form of the function fitted to these points.
@@ II.2 Standard Zero-Noise Extrapolation
In particular, this is true if the noisy implementations ${\mathcal{E}}_{U}$ and ${\mathcal{E}}_{U^{\dagger}}$ of both, $U$ and $U^{\dagger}$, amount to a depolarising error $\Lambda_{p}$ with error probability $p$, i.e.
@@
with maximally mixed state $\rho_{*}=I/2^{q}$ (where $I$ denotes the identity and $q$ the number of qubits).
In this case, the noisy implementation of the sequence $U(U^{\dagger}U)^{n}$ indeed yields again a depolarising error with scaled error probability $\lambda p$ (in first order of $p$) [9, 11]. This can be traced back to the fact that depolarising errors commute with unitary operations, i.e., $\Lambda_{p}{\mathcal{U}}={\mathcal{U}}\Lambda_{p}$ and $\Lambda_{p}{\mathcal{U}}^{\dagger}={\mathcal{U}}^{\dagger}\Lambda_{p}$, and hence
@@
$\simeq$ $\Lambda_{(2n+1)p}{\mathcal{U}}$
for small $p$. Indeed, we obtain the same result as in Eq. (6) with re-scaled probability of the depolarising error ($p\to\lambda p$, where $\lambda=2n+1$). For other error models, however, a similarly simple scaling behaviour does in general not hold. In the case of coherent errors, it may even happen that the errors of ${\mathcal{E}}_{U}$ and ${\mathcal{E}}_{U^{\dagger}}$ exactly compensate each other, such that ${\mathcal{E}}_{U^{\dagger}}{\mathcal{E}}_{U}=I$ and the error is not amplified at all.
Due to its simplicity and moderate sampling overhead, the method of ZNE based on identity insertions is frequently used in order to mitigate gate errors. In this article, we will use it as a benchmark for our more refined method introduced in Sec. III. More specifically, we employ the following version, subsequently referred to as ‘standard ZNE’ (sZNE): first, we concentrate on two-qubit gate (i.e., CNOT) errors, which are the dominant source of errors in the superconducting device provided by IBM (i.e., one or two orders of magnitude larger than the errors of single-qubit gates) [20, 21, 22, 23, 24]. We use noise scaling factors $\lambda=1$ (the original circuit), $\lambda=3$ (every CNOT gate replaced by 3 CNOT gates) and $\lambda=5$ (5 CNOT gates) and fit the measured data points by a linear function. In order to evaluate the expectation value $\langle A\rangle$, we employ readout error mitigation based on the method M3 [25].
II.3 Randomised Compiling
Besides readout error mitigation, additional techniques can be used to enhance the performance of ZNE [26, 6, 14, 27, 28, 29]. In this paper, we will investigate the effect of randomised compiling or so called Pauli twirling [30] on ZNE. The technique of Pauli twirling is based on the fact that the CNOT gate is a Clifford gate [31], and hence every Pauli operator is transformed again into a Pauli operator after sandwiching it between two CNOT operations:
${\rm CNOT}\leavevmode\nobreak\ {\mathcal{P}}_{i}\leavevmode\nobreak\ {\rm CNOT}={\mathcal{P}}_{\pi(i)},$ (10)
@@ II.3 Randomised Compiling
${\mathcal{P}}_{i}\leavevmode\nobreak\ {\rm CNOT}\leavevmode\nobreak\ {\mathcal{P}}_{\pi(i)}={\rm CNOT}.$ (11)
Sandwiching each CNOT gate between two Pauli operators ${\mathcal{P}}_{i}$ and ${\mathcal{P}}_{\pi(i)}$ with randomly chosen $i$ hence preserves the CNOT logic in absence of noise. In presence of noise, however, the error channel $\Lambda$ of the noisy CNOT gate ${\mathcal{E}}_{\rm CNOT}=\Lambda\leavevmode\nobreak\ {\rm CNOT}$ is transformed, on average, into a Pauli channel $\Lambda_{\mathcal{P}}$ of the following form:
@@
In other words, the Pauli operator ${\mathcal{P}}_{i}$ is applied with probability $p_{i}$, and thus produces an error if ${\mathcal{P}}_{i}\neq I$. Twirling is also applied to adjacent qubits, which could suffer from crosstalk [32, 33], as described in [34]. Each of the twirling gates is combined with other single-qubit operations to avoid increasing circuit length and therefore time (see Fig. 1). Note that if the noisy CNOT gate originally suffers from a coherent error (i.e., if its error channel $\Lambda$ is unitary), the latter is transformed into an incoherent error by Pauli twirling. As already hinted at in Sec. II.2, the scaling behaviour of the noise due to repeated CNOT gates is expected to be more accurate in this case.
Figure 1: On the left: Example of a bare circuit with single qubit gates and one CNOT gate. In the middle: Twirling is performed by adding random Pauli gates (rectangles with hatching) to the circuit without changing the logic of the CNOT gates in the circuit. Twirling is only applied before and after CNOT gates, as well as to adjacent qubits, which could be affected by the CNOT gate due to crosstalk. Finally, the number of single qubit gates is reduced by contraction of subsequent single-qubit gates (circuit one the right).
@@ III Inverted-Circuit Zero-Noise Extrapolation
As explained in Sec. II.2, the standard ZNE relies on a certain assumption concerning the scaling of the error strength, namely that the error of a gate is amplified by a factor $\lambda$ if the gate is repeated $\lambda-1$ times. In general, however, this is only true for small errors commuting with the ideal gate operation (such as depolarising errors). We therefore propose not to rely on this assumption and, instead, to measure the error strength $\epsilon$ of a quantum circuit $U$ used to prepare state $|\psi\rangle$ directly.
@@ III.1 Definition of error strength $\epsilon$
The operator $\sigma$ represents the noisy part. In case of depolarising errors, we obtain $\sigma=(I-|\psi\rangle\langle\psi|)/(2^{q}-1)$ (independent of $\epsilon$) and the following relation between the noise strength $\epsilon$ and the error probability $p$ of the depolarising channel: $\epsilon=p\left(1-\frac{1}{2^{q}}\right)$. In general, however, $\sigma$ and $\epsilon$ are, a priori, unknown. Note that $\sigma$ is not even guaranteed to be positive, but this poses no problem in the following (where we will not assume that $\sigma$ is positive). From Eqs. (13-15), it follows that:
@@
The expectation value of an observable $A$ then can be written as follows:
$\langle A\rangle=(1-\epsilon)\underbrace{\langle\psi|A|\psi\rangle}_{=\langle A\rangle_{\rm ideal}}+\epsilon\leavevmode\nobreak\ {\rm tr}(A\sigma).$ (17)
It converges to $\langle A\rangle_{\rm ideal}$, see Eq. (1), if $\epsilon\to 0$. Moreover, we see that a linear fit of $\langle A\rangle$ as a function of $\epsilon$ should be adequate for performing ZNE, at least if the dependence of the noisy part $\sigma$ on $\epsilon$ is negligible. Regarding the latter point, the technique of randomised compiling may be useful (refer to Sec.  II.3) as it transforms coherent errors into incoherent ones.

### v3

@@ Abstract
A common approach to deal with gate errors in modern quantum-computing hardware is zero-noise extrapolation. By artificially amplifying errors and extrapolating the expectation values obtained with different error strengths towards the zero-error (zero-noise) limit, the technique aims at rectifying errors in noisy quantum computing systems. For accurate extrapolation, it is essential to know the exact factors of the noise amplification. In this article, we propose a simple method for estimating the strength of errors occurring in a quantum circuit and demonstrate improved extrapolation results. The method determines the error strength for a circuit by appending to it the inverted circuit and measuring the probability of the initial state. The estimation of error strengths is easy to implement for arbitrary circuits and does not require previous characterization of noise properties. We compare this method with the conventional zero-noise extrapolation method and show that our method leads to a more accurate calculation of expectation values on current quantum hardware, showcasing its suitability for near-term quantum computing applications.
@@ I Introduction
Current quantum computing devices, so-called Noisy Intermediate Scale Quantum (NISQ) hardware [1], suffer from errors [1, 2, 3, 4]. The accuracy of the output is reduced by these errors. Quantum error mitigation techniques aim to minimize these inaccuracies. The general principle of these techniques is to deduce an almost noise-free estimate of the expectation value of interest from results obtained on noisy hardware. For example, the technique of zero-noise extrapolation (ZNE) [2, 5, 6, 3, 7, 8] consists of performing a series of measurements with systematically amplified errors. The expectation values obtained with different error strengths are then extrapolated to zero error (zero-noise). Hence, a more accurate estimation of the expectation value can be obtained at the cost of an increased sampling overhead due to a larger number of circuits and classical post-processing.
@@
Several zero-noise extrapolation schemes have been proposed, each with their own advantages and challenges. In continuous ZNE [2, 5], the microwave pulses generating the elementary single- and two-qubit gates are scaled to increase the noise throughout the circuit. This extrapolation yields satisfactory results, however, it requires a time-consuming calibration of the scaled pulses. The implementation of the discrete ZNE method [9, 10, 11, 12] is more straightforward. Here, sequences of gates are inserted that, in the error-free case, are equivalent to the identity. Depending on the details of the implementation, different noise scaling factors can be targeted; for example, replacing every CNOT gate with three CNOT gates corresponds to a noise-scaling factor of three. While this simple scaling behavior holds, for example, for depolarizing errors to the limit of small error strength, this is usually not the case for realistic noise models. Depending on the specific type of errors occurring on the given hardware, not every error is amplified exactly $\lambda$ times if the corresponding gate is executed $\lambda$ times [13, 14, 15]. This may lead to a certain bias concerning the estimation of the error strength, which deteriorates the results of the discrete ZNE method [13, 14]. Another powerful approach is ZNE with Probabilistic Error Amplification (PEA) [16]. With PEA, a noise model is first learned for different circuit layers, and errors are then amplified by inserting random gates according to the noise model. This approach has a high sampling cost due to the larger number of experiments required to learn the noise model. Probabilistic Error Cancellation (PEC) [2, 17] is another method with an even higher sampling cost [18] that represents the ideal circuit as a quasi-probability distribution of previously characterized noisy ones. Finally, if the noise can be described by a depolarizing noise model, the expectation value can be corrected after measuring the strength of the depolarizing noise [19, 20] using a method based on inverted circuits similar to the one which we will introduce below in a more general context. This correction is easy to implement, but it is only applicable in the case of errors well approximated by depolarizing noise.
@@
In this article, by building on the relatively simple method of ZNE with identity insertions [9], we introduce the method of Inverted-Circuit-ZNE (IC-ZNE) which allows for the measurement of the entire error of a circuit to obtain an exact noise scaling parameter for better error mitigation. This involves adding an extra circuit with its inverse to determine the actual error strength $\epsilon$. This method overcomes the above problem that not every error type is amplified $\lambda$ times when increasing the number of noisy gates by $\lambda$ [6]. We further explicitly show that the total error is actually different from the assumed error amplification. Note that our IC-ZNE approach does not require any knowledge of the underlying quantum computer noise model.
Methods that also involve an inverted circuit are (dual-state) purification [21, 22] and verified phase estimation [23]. The main idea of dual-state purification is to consider $\rho^{2}/{\rm}Tr(\rho^{2})$ as a purified version of the noisy quantum state $\rho$. When combined with ZNE (as in Refs. [24, 25]), this method can be regarded as complementary to our IC-ZNE approach: concerning the expectation values obtained for different circuits with scaled noise, dual-state purification yields more precise (purified) expectation values, whereas our method focuses on the accurate determination of the corresponding noise strengths.
A more general framework (containing the above-mentioned dual-state purification as sub-version) is provided by the generalized quantum subspace expansion [24, 25]. Another sub-version (fault subspaces) is related to ZNE, since, similar as in ZNE, it considers superpositions of states created by circuits with different noise strengths. In contrast to ZNE, the coefficients of this superposition are obtained from a variational principle and not by extrapolation to zero noise. This method can be applied to variational problems (e.g., finding the ground state energy of a given Hamiltonian), where it appears to be well suited due to its robustness against inaccurate knowledge of noise levels [25]. In contrast, our IC-ZNE approach can be used to mitigate expectation values for arbitrary observables and quantum circuits, not restricted to variational problems.
Finally, the verified phase estimation approach mentioned above uses inverted circuits to detect and discard errors [23]. However, it can be applied only for the purpose of phase estimation (or more generally, if the prepared quantum state is an eigenstate of the measured observable).
@@
The paper is organized as follows. After introducing known concepts and methods used in this paper like standard ZNE (sZNE) and randomized compiling in Sec. II, we introduce the method of Inverted-Circuit ZNE (IC-ZNE) in Sec. III. Starting from a precise definition of the error strength $\epsilon$, we derive a relation between $\epsilon$ and the probability of measuring all qubits in their initial state after the inverse circuit has been added. Finally, we demonstrate and explain the improved performance of our method compared to standard ZNE on simulators with depolarizing noise and on IBM’s quantum computing hardware in Sec. IV. A conclusion is drawn in Sec. V.
@@ II.1 General setting and notation
The main idea of ZNE now consists of introducing a parameter $\epsilon$ which is expected to describe the amount of noise present in ${\mathcal{E}}_{U}$. Assuming that this parameter, or at least its scaling $\lambda=\epsilon/\epsilon_{0}$ with respect to the (in general unknown) noise strength $\epsilon_{0}$ of the original circuit, can be controlled by adding additional noise to the system, the data points $\langle A\rangle_{\lambda_{i}}$ measured at different noise scaling factors $\lambda_{i}$ can be extrapolated to yield $\langle A\rangle_{\rm ideal}=\langle A\rangle_{\lambda\to 0}$. Different versions of ZNE differ in the way in which the noise parameter $\epsilon$ (or, equivalently, the noise scaling factor $\lambda$) is defined, the number of data points to be measured at different values of the noise strengths and the form of the function fitted to these points.
@@ II.2 Standard Zero-Noise Extrapolation
In particular, this is true if the noisy implementations ${\mathcal{E}}_{U}$ and ${\mathcal{E}}_{U^{\dagger}}$ of both, $U$ and $U^{\dagger}$, amount to a depolarizing error $\Lambda_{p}$ with error probability $p$, i.e.
@@
with maximally mixed state $\rho_{*}=I/2^{q}$ [26] (where $I$ denotes the identity and $q$ the number of qubits).
In this case, the noisy implementation of the sequence $U(U^{\dagger}U)^{n}$ indeed yields again a depolarizing error with scaled error probability $\lambda p$ (in first order of $p$) [9, 11]. This can be traced back to the fact that depolarizing errors commute with unitary operations, i.e., $\Lambda_{p}{\mathcal{U}}={\mathcal{U}}\Lambda_{p}$ and $\Lambda_{p}{\mathcal{U}}^{\dagger}={\mathcal{U}}^{\dagger}\Lambda_{p}$, and hence
@@
$=$ ${\Lambda_{1-(1-p)^{2n+1}}~{}{\mathcal{U}}}.$
Indeed, we obtain the same result as in Eq. (6) with re-scaled probability of the depolarizing error, i.e., $1-p\to(1-p)^{\lambda}$, where $\lambda=2n+1$, which simplifies to $p\to\lambda p$ for $\lambda p\ll 1$. The expectation value of $A$ results as:
$\langle A\rangle=(1-p)^{\lambda}\langle A\rangle_{\rm ideal}+\left[1-(1-p)^{\lambda}\right]\langle A\rangle_{*},$ (10)
where $\langle A\rangle_{*}={\rm tr}(A\rho_{*})$, or
$\langle A\rangle\simeq(1-\lambda p)\langle A\rangle_{\rm ideal}+\lambda p\langle A\rangle_{*}$ (11)
for $\lambda p\ll 1$.
For other error models, however, a similarly simple scaling behavior generally does not hold. In the case of coherent errors, it may even happen that the errors of ${\mathcal{E}}_{U}$ and ${\mathcal{E}}_{U^{\dagger}}$ exactly compensate for each other, such that ${\mathcal{E}}_{U^{\dagger}}{\mathcal{E}}_{U}=I$ and the error is not amplified at all.
Due to its simplicity and moderate sampling overhead, the method of ZNE based on identity insertions is frequently used in order to mitigate gate errors. In this article, we will use it as a benchmark for our more refined method introduced in Sec. III. More specifically, we employ the following version, subsequently referred to as ‘standard ZNE’ (sZNE): first, we concentrate on two-qubit gate (i.e., CNOT) errors, which are the dominant source of errors in the superconducting device provided by IBM (i.e., one or two orders of magnitude larger than the errors of single-qubit gates) [27, 28, 29, 30, 31]. We use noise scaling factors $\lambda=1$ (the original circuit), $\lambda=3$ (every CNOT gate replaced by 3 CNOT gates) and $\lambda=5$ (5 CNOT gates) and fit the measured data points by an exponential function $f(\lambda)=a_{1}e^{-a_{2}\lambda}+a_{3}$ motivated by Eq. (10), which is generally expected to yield more accurate results than a linear fit based on Eq. (11) [32]. When performing the exponential fit, we restrict the parameters $a_{1}$ and $a_{3}$ to the interval $[A_{\rm min},A_{\rm max}]$ given by the minimum and maximum eigenvalue of $A$. Without this restriction, we have observed that the exponential fit may be highly unstable in some cases, especially if noise strengths are relatively large. Furthermore, in order to evaluate the expectation value $\langle A\rangle$, we employ readout error mitigation based on the method M3 [33].
II.3 Randomized compiling
Besides readout error mitigation, additional techniques can be used to enhance the performance of ZNE [34, 6, 14, 35, 36, 37]. In this paper, we will investigate the effect of randomized compiling or so-called Pauli twirling [38] on ZNE. The technique of Pauli twirling is based on the fact that the CNOT gate is a Clifford gate [39], and hence every Pauli operator is transformed into a Pauli operator again, after sandwiching it between two CNOT operations:
${\rm CNOT}~{}{\mathcal{P}}_{i}~{}{\rm CNOT}={\mathcal{P}}_{\pi(i)},$ (12)
@@ II.3 Randomized compiling
${\mathcal{P}}_{i}~{}{\rm CNOT}~{}{\mathcal{P}}_{\pi(i)}={\rm CNOT}.$ (13)
Sandwiching each CNOT gate between two Pauli operators ${\mathcal{P}}_{i}$ and ${\mathcal{P}}_{\pi(i)}$ with randomly chosen $i$ hence preserves the CNOT logic in the absence of noise. In the presence of noise, however, the error channel $\Lambda$ of the noisy CNOT gate ${\mathcal{E}}_{\rm CNOT}=\Lambda~{}{\rm CNOT}$ is transformed, on average, into a Pauli channel $\Lambda_{\mathcal{P}}$ of the following form:
@@
In other words, the Pauli operator ${\mathcal{P}}_{i}$ is applied with probability $p_{i}$, and thus produces an error if ${\mathcal{P}}_{i}\neq I$. Twirling is also applied to adjacent qubits, which could suffer from crosstalk [40, 41], as described in [42]. Each of the twirling gates is combined with other single-qubit operations to avoid increasing circuit length and therefore time (see Fig. 1). Note that if the noisy CNOT gate originally suffers from a coherent error (i.e., if its error channel $\Lambda$ is unitary), the latter is transformed into an incoherent error by Pauli twirling. As already hinted at in Sec. II.2, the scaling behavior of the noise due to repeated CNOT gates is expected to be more accurate in this case.
Figure 1: On the left: Example of a bare circuit with single qubit gates and one CNOT gate. In the middle: Twirling is performed by adding random Pauli gates (rectangles with hatching) to the circuit without changing the logic of the CNOT gates in the circuit. Twirling is only applied before and after CNOT gates, as well as to adjacent qubits, which could be affected by the CNOT gate due to crosstalk. Finally, the number of single qubit gates is reduced by contraction of subsequent single-qubit gates (circuit on the right).
@@ III Inverted-Circuit Zero-Noise Extrapolation
As explained in Sec. II.2, the standard ZNE relies on a certain assumption concerning the scaling of the error strength, namely that the error of a gate is amplified by a factor $\lambda$ if the gate is repeated $\lambda-1$ times. In general, however, this is only true for small errors commuting with the ideal gate operation (such as depolarizing errors). We therefore propose not to rely on this assumption and, instead, to measure the error strength $\epsilon$ of a quantum circuit $U$ used to prepare state $|\psi\rangle$ directly.
@@ III.1 Definition of error strength $\epsilon$
The operator $\sigma$ represents the noisy part. In case of depolarizing errors, we obtain $\sigma=(I-|\psi\rangle\langle\psi|)/(2^{q}-1)$ (independent of $\epsilon$) and the following relation between the noise strength $\epsilon$ and the error probability $p$ of the depolarizing channel: $\epsilon=p\left(1-\frac{1}{2^{q}}\right)$. In general, however, $\sigma$ and $\epsilon$ are, a priori, unknown. Note that $\sigma$ is not even guaranteed to be positive, but this poses no problem in the following (where we will not assume that $\sigma$ is positive). From Eqs. (15-17), it follows that:
@@
The expectation value of an observable $A$ can then be written as follows:
$\langle A\rangle=(1-\epsilon)\underbrace{\langle\psi|A|\psi\rangle}_{=\langle A\rangle_{\rm ideal}}+\epsilon~{}{\rm tr}(A\sigma).$ (19)
It converges to $\langle A\rangle_{\rm ideal}$, see Eq. (1), for $\epsilon\to 0$. Moreover, note that, due to our definition of $\epsilon$ based on the fidelity, cf. Eq. (16), the error strength $\epsilon$ is restricted to the interval $[0,1]$ – in contrast to the noise scaling factor $\lambda$ based on the number of gate repetitions, which may, in principle, be chosen arbitrarily large. For this reason, a linear fit of $\langle A\rangle$ as a function of $\epsilon$ should be sufficiently accurate for IC-ZNE, as is evident from Eq. (19), at least if the dependence of the noisy part $\sigma$ on $\epsilon$ is negligible. Regarding the latter point, the technique of randomized compiling may be useful (refer to Sec.  II.3) as it transforms coherent errors into incoherent ones.